OpenStudy (anonymous):

the area of a circle is given by (3.14k^2-12*3.14k+36*3.14) square inches. What is the radius? I substituted the pi sign for 3.14 lol

OpenStudy (anonymous):

\[\pi*k ^{2} - 12\pi*k + 36\pi\] \[\pi(k^2 - 12k + 36)\] \[\pi(k - 6)^2\] Formula of Circle is \[\pi*r^2\] so raidus is just (k - 6)